OpenStudy (anonymous):

Solve for x. 7 - (2x + 11) + 3(3 - x) = 20

OpenStudy (anonymous):

I no longer need help on this question, but can you help me with this: Given: x - 7 > -2. Choose the solution set. {x | x R, x > -9} {x | x R, x > -5} {x | x R, x > 5} {x | x R, x > 14}

OpenStudy (redohawk):

7 - (2x + 11) + 3(3 - x) = 20 -14x - 77 + 33 - 3x = 20 -17x - 44 = 30 Stop! you don't need help anymore? kk

OpenStudy (anonymous):

I don't know what that question is even asking. derp.

OpenStudy (redohawk):

why is their an R in the answer but not in the question. no sense whatsoever... I am lost !!!

OpenStudy (anonymous):

I think R means Real Number